OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a black marble, replacing it, and then picking a blue marble? STEPS ON IT?

OpenStudy (anonymous):

there are a total of 15 marbles in the bag, so that will be your denominator

OpenStudy (anonymous):

the probability that the first marble selected is black is the ratio of the black marbles to the total, i.e. \[\frac{4}{15}\]

OpenStudy (anonymous):

then the probability that the second marble chosen is blue is also the ratio of the number of blue marbles to the total number of marbles, i.e. \[\frac{6}{15}=\frac{2}{5}\]

OpenStudy (anonymous):

the probability that BOTH things occur, i.e. the first is black AND the second is blue is the product of those to numbers. namely \[\frac{4}{15}\times \frac{2}{5}\]

OpenStudy (anonymous):

*those two numbers

OpenStudy (anonymous):

hope it is clear, all steps are there

OpenStudy (anonymous):

yw